A roller-skater has a difficult trick with a 16% of success. They try the trick 25 times and want to know the probability of exactly 4 successes.

Compute:

(a) the exact answer

(b) the Poisson approximation

(c) the normal approximation (Hint: you must use the histogram correction.)
Answer #1

Solhon :- þ= 0.16 m = 25 25-4 a) Exact, xn Binomial P(x) = ( ) pt 9- P(x= 4) = (25) (0.16)* (0.845 = 0.2130 b) a= mp = 25x0.1